A local hamburger shop sold a combined total of 588 hamburgers and cheeseburgers on Tuesday. There were 62 fewer cheeseburgers sold than hamburgers. How many hamburgers were sold on Tuesday?
Answer by mananth(16949)

cheeseburgers ------- x numbers
hamburgers----------- y numbers
y-x= 62
x+y = 588
add both equations
2y=650
y=650/2
y= 325 ---------- hamburgers
cheese burgers 325-62= 263